Explain the equation of parabola in detail with examples

Grade:12th pass

1 Answers

utpal raj
7 years ago
1) Describe the curve represented by 2x2 - x - 4y + 7 = 0.
Solution: - (1/4)x + (3/4)This is a parabola which opens upward. The vertex is located where x = (-1/4)/(2[1/2]) = 1/4. If x = 1/4, then y = 23/32. Thus, the vertex is at (1/4, 23/32). 2 = x - 3y = (1/2)x2 - x - 4y + 3 = 0-4y = -2x22x